a campground rents campsites for $12 per night. at this rate, all 90 campsites are usually rented. for each $1 increase in the price per night, about 3 less sites are rented. the campground's nightly revenue can be modeled by R=(90-3x)(12+x). use the vertex form to find how the campground can maximize nightly revenue
Answer by stanbon(75887)

R=(90-3x)(12+x). use the vertex form to find how the campground can maximize nightly revenue
----------
R(x) = -3x^2-36x+90x+12*90
R(x) = -3x^2+ 54x + 1080
-----------------
Convert to vertex form by completing the square:
R(x)-1080-3*? = -3(x^2-18x+?)
R(x)-1080-3*81 = -3(x^2-18x+81)
R(x)-1323 = -3(x-9)^2
---
Vertex: (9,1323)
Increase the price by $9; Revenue will be $1323.
Cheers,
Stan H.